OpenStudy (anonymous):

Solve for a: 2^a/ln2-2a=1/ln2

OpenStudy (kc_kennylau):

\[\frac{2^a}{\ln2}-2a=\frac1{\ln2}\]

OpenStudy (kc_kennylau):

Combine the first two terms first

OpenStudy (anonymous):

How?

OpenStudy (kc_kennylau):

Multiply the second term by \(\frac{\ln2}{\ln2}\)

OpenStudy (anonymous):

so (2^a-lna(a))/ln2

OpenStudy (anonymous):

sorry- ln2 instead of lna

OpenStudy (kc_kennylau):

Yes

OpenStudy (kc_kennylau):

Then combine the first two terms

OpenStudy (kc_kennylau):

Actually you can multiply the whole thing by \(\ln2\)

OpenStudy (kc_kennylau):

So multiply the whole thing by ln2

OpenStudy (anonymous):

so \[2^a-\ln2(a)=1\]

OpenStudy (kc_kennylau):

Yep

OpenStudy (anonymous):

make that \[2^a-2\ln2(a)=1\]

OpenStudy (anonymous):

i forgot the 2..

OpenStudy (kc_kennylau):

oh I didn't notice too sorry :P

OpenStudy (anonymous):

now what?

OpenStudy (kc_kennylau):

Would you please make the a in front of the ln 2

OpenStudy (anonymous):

sure: \[2^a-2a(\ln2)=1\]

OpenStudy (kc_kennylau):

Well sorry I need to think for a while :P

OpenStudy (anonymous):

that's what happened to me.. I'm really clueless when it comes to this stuff.. transcendentals are not my thing.

OpenStudy (kc_kennylau):

It's not really solvable, double-check your question perhaps?

OpenStudy (anonymous):

no, it's the right equation... can you perhaps complete the square with it?

OpenStudy (kc_kennylau):

One solution is a=0

OpenStudy (kc_kennylau):

The another is a = (-2 W_(-1)(-1/(2 sqrt(e)))-1)/(2 log(2))

OpenStudy (kc_kennylau):

I've never met the function W

OpenStudy (anonymous):

Calculus fail.. basically just plug one side into the calculator and the other side and see where they intersect?

OpenStudy (kc_kennylau):

Please triple-check your question for any missing parentheses

OpenStudy (anonymous):

do you want the original?

OpenStudy (kc_kennylau):

Yes

OpenStudy (anonymous):

Find a value a such that the average value of the function f(x)=\[2^x-1\] on the interval [0,a] is equal to one.

OpenStudy (kc_kennylau):

That means \(\Large\displaystyle\int^a_02^x-1dx=1\)?

OpenStudy (anonymous):

but the integral must be divided by a as that is the rule for averaage value- area divided by length of interval gives average value

OpenStudy (kc_kennylau):

oh sorry

OpenStudy (kc_kennylau):

That means \(\displaystyle\Large\frac1a\int^a_0(2^x-1)dx=1\)?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

and the anti-derivative should be \[(2^x)/\ln2-x\]

OpenStudy (kc_kennylau):

2^0 is not 0 !!!

OpenStudy (anonymous):

I know, it's 1

OpenStudy (anonymous):

therefore, you get -(1/ln2)

OpenStudy (kc_kennylau):

Where did the 1 go?

OpenStudy (anonymous):

not sure..good catch

OpenStudy (anonymous):

wait, it's still there..

OpenStudy (kc_kennylau):

2^a/ln2-2a=1/ln2+1

OpenStudy (anonymous):

not +1 because x=0

OpenStudy (kc_kennylau):

The average is 1

OpenStudy (anonymous):

By the fundamental theorem of calculus, that x should change to 0, because 0 is the lower limit of integration.

OpenStudy (kc_kennylau):

(2^a/ln2-2a)-(1/ln2-0) = 1

OpenStudy (anonymous):

yes, and back to where we started... unfortunately...

OpenStudy (kc_kennylau):

Therefore 2^a/ln2-2a=1/ln2+1

OpenStudy (anonymous):

remember that we have to divide the entire LH side by a... and that should be just a, not 2a, by the FTC

OpenStudy (kc_kennylau):

Actually I'm sick so my brain is not functioning properly

OpenStudy (kc_kennylau):

So 2^a/ln2-a=1/ln2+1

OpenStudy (anonymous):

change the 1 on the RH side to a and you got it...

OpenStudy (kc_kennylau):

Wait where did the 1/a go

OpenStudy (anonymous):

there never was one.. the integral needed to be divided by a, and then you multiply both side by a to get it out of the denominator, changing the RH side to 1(a), or a

OpenStudy (kc_kennylau):

Oh

OpenStudy (kc_kennylau):

So your original question was correct sorry

OpenStudy (anonymous):

yeah.. I tihnk I found a way to do it on the calculator though.... I believe you must somehow use quadratics..